OpenStudy (anonymous):

Help please re simplifying a log expression 5xlog base 3 (1/x) + log base 3 (x^2) - 2xlog base 3 (sqrt x^3)

OpenStudy (anonymous):

Please :)

OpenStudy (anonymous):

log x + log y= log (x*y) log x - log y = log (x/y) x log y = log y^x

OpenStudy (anonymous):

So is it then 1/x * x^2 then divide by x^1/2

OpenStudy (anonymous):

not quite right

OpenStudy (anonymous):

there is 5x

OpenStudy (anonymous):

5x log_3 (1/x)= log_3(1/x)^5x

OpenStudy (anonymous):

log_3((1/x)^5x * x^2)

OpenStudy (anonymous):

but doesn't that then give me a negative?

OpenStudy (anonymous):

log_3 (x^2/x^5)= log_3 (1/x^3)=log_3 (1/x)^3= 3 log_x (1/x)

OpenStudy (anonymous):

now you have 3log_3(1/x)- 2x log_3 (sqrt(x^3))

OpenStudy (anonymous):

now just do the same thing for the last part

OpenStudy (anonymous):

is that then 1/x^3 divided by x^3/2?

OpenStudy (anonymous):

log_3 ((1/x^3)/(sqrt(x^3))^5x)

OpenStudy (anonymous):

log_3 ((1/x^3)/((x^3/2)^5x)= log_3((1/x^3)/(x^15x/2) log_3(1/(x^3*x^15x/2) log_3(1/x^(15x/2+3) ...

OpenStudy (anonymous):

Sorry I am an idiot, why raise x^3/2 to ^5?

OpenStudy (anonymous):

Thank you for all of your help xoxo